jake has a motor bike. he travels to and from work three times in a week, he also travels to and from his girlfriends house, which is one-third of the distance to work four times a week. in total he travels 156 km. how far is it from his house to work?
Answer by mananth(16946)

distance to work =x
3 times = 3x
distance to girlfriends house =x/3
4 times = 4x/3
total distance = 156
3x+(4x/3) =156
X3
9x+4x=3*156
13x=3*156
x= 36
distance to work 36km